# Decoding Decoded: Understanding Hyperparameter Effects in Open-Ended Text Generation

### v1

@@ 6 Practical Recommendations
Based on the results of our study, we provide the following practical recommendations for selecting decoding strategies in open-ended text generation tasks. To achieve a balance between coherence and diversity, we recommend using contrastive search with moderate values of $\alpha$ and $k$. Specifically, setting $\alpha$ to 0.6 and $k$ to either 5 or 10 has consistently led to high-quality text outputs in our experiments. These parameter ranges strike an effective balance between generating coherent and diverse text, as visualized in the appendix. Our analysis shows a sharp performance decline when moving toward extreme values, such as low $\alpha$ or low $k$ (which lead to overly coherent but dull text), or high $\alpha$ and high $k$ (which produce diverse but incoherent text).
In addition to manually tuned hyperparameters, we recommend considering hyperparameter-free solutions, such as adaptive contrastive search, which dynamically adjusts the parameters during generation. This approach mitigates the need for fine-tuning specific values and has shown promising results in generating well-balanced text without requiring extensive parameter exploration.
On the other hand, beam search is generally not recommended for open-ended text generation. While beam search often excels in tasks that require high coherence, it tends to produce repetitive and less diverse text, making it unsuitable for creative or varied content generation.
Sampling methods, including top-k and nucleus (p) sampling, can perform well but tend to exhibit higher variance in output quality. Among these methods, we recommend using top-k sampling with $k=50$ and nucleus sampling with $p=0.9$ or $p=0.95$. These parameter settings have consistently produced high-quality text that aligns well with both automatic and human evaluation metrics. While these strategies are effective, their inherent randomness can sometimes lead to variability, so multiple generations may be necessary to achieve optimal results.
@@
It is also important to caution against relying solely on automatic evaluation metrics when assessing text quality. Our results demonstrate that current automatic metrics are insufficient to fully capture human preferences. As noted in previous studies, including Garces-Arias et al. (2024), certain decoding methods, such as contrastive search with the DoubleExp method, achieve high scores on metrics like MAUVE but are consistently rejected by human evaluators for lack of coherence or clarity. This suggests that more research is needed to develop robust evaluation metrics that better correlate with human judgments of text quality.
Lastly, we encourage users to consider the specific dataset when selecting decoding strategies. For example, tasks like story generation (as in the book prompt) tend to reward creativity, so strategies that promote diversity may be preferable. In contrast, tasks like Wikipedia or news generation (as in Wikinews) demand more factual accuracy and coherence, where stricter control over diversity may be more beneficial. Detailed performance analyses for these different scenarios are provided in the appendix, offering further insights into which strategies work best for each dataset type.
@@ 7 Conclusion
Decoding strategies for large language models (LLMs) are a crucial yet often underexamined aspect of open-ended text generation. This study presents a detailed analysis of how hyperparameter selection across various decoding strategies significantly influences the quality of generated text. The findings highlight the necessity of maintaining a balance between coherence and diversity, as strategies that heavily prioritize one tend to underperform overall. Through extensive sensitivity analysis, we show that the choice of decoding method and its associated hyperparameters can impact text quality as much as, if not more than, model size.
@@
While this study provides insights into how hyperparameters affect the quality of generated text, it also has limitations. First, the study is confined to open-ended text generation tasks, and further research is necessary to assess the applicability of our methods to other NLP tasks such as summarization and machine translation. Additionally, the impact of our findings on models that have undergone supervised fine-tuning (SFT) remains unexplored. Future work could investigate how these models perform under similar experimental conditions.
Moreover, the analysis was conducted exclusively in English, leaving the generalizability to multilingual settings untested. Previous studies Su and Collier (2023); Garces-Arias et al. (2024) suggest that these findings may generalize to other languages, but further investigation is required to confirm this.
A key observation from our experiments is the discrepancy between individual automatic evaluation metrics and human judgments, highlighting the need for more reliable automatic evaluation methods. Given the resource-intensive nature of human evaluations, developing comprehensive automatic metrics that more accurately capture both coherence and diversity, while better aligning with human preferences, remains a critical area for future work.
Due to resource constraints, conducting an extensive human evaluation covering more combinations of models and decoding strategies was not feasible. This limitation underscores the necessity for automatic metrics that better correlate with human judgment, as a comprehensive human evaluation would be prohibitively costly.
Finally, this study focuses on analyzing the sensitivity of commonly used automatic metrics, which are inherently influenced by the models employed in the evaluation process. For example, coherence is assessed using the OPT 2.7B model, and MAUVE is calculated with GPT-2-based embeddings. The potential biases introduced by these models have not been fully explored, and addressing these biases in future research may offer valuable insights.
@@ Acknowledgments
Matthias Aßenmacher received funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of BERD@NFDI, under grant number 460037581.
Appendix
Appendix A1 Case Studies
@@
Appendix A2 Performance per Model, Dataset and Decoding Strategy
@@
Figure 5: Effect of beam width on metric behavior. Falcon2 is used for visualization, as the trends observed were consistent across all models.
@@
Appendix A3 Correlation of Human evaluation and Automatic metrics
Figure 26: Correlation between human scores and automatic metrics. Moderate correlations are observed between human judgments and automatic metrics such as QText and MAUVE, while Coherence and Diversity show weak to very weak correlations.
Appendix A4 Human Evaluation Form

### v2

@@ 6 Practical Recommendations
Based on the results of our study, we provide the following practical recommendations for selecting decoding strategies in open-ended text generation.
Deterministic Strategies
Beam search is generally not recommended for open-ended text generation tasks due to its propensity to produce outputs that fail to capture the diversity of human language. As illustrated in Figure 4, our findings demonstrate consistently low performance across all models and beam width settings when compared to human references across four evaluation metrics. A notable trend is that smaller beam widths (e.g. 3 or 5) tend to yield better results; however, even at these settings, the performance disparity remains substantial, particularly concerning MAUVE scores and diversity measures. While the Falcon2 (11B) model emerged as the best performer in this task, it still underperforms across all four metrics.
Sampling Strategies
Sampling with temperature is highly sensitive to its temperature hyperparameter. Figure 20 illustrates that performance increases monotonically for all models and metrics as temperature rises in $t\in[0.1,1.0]$. The best results are consistently observed at temperatures of 0.9 or 1.0. Both Falcon2 and Qwen2 perform best overall, though all models achieve strong results at higher temperatures. Larger models do not show a clear advantage over smaller ones. Specifically, GPT2-XL (1.5B) might outperform Llama3 (8B) under the same hyperparameter configuration.
For top-$k$ sampling, performance is also sensitive to the truncation hyperparameter $k$. As shown in Figure 22, small $k$ values (e.g., $k=3,5$) produce lower performance, while medium to larger $k$ values (e.g., $k=10,15,20,50$) yield higher performance. The highest results occur at $k=20$ or $k=50$, and performance similar to human references is achievable with $k=10$ or $k=15$. Qwen2 reaches the best overall performance, but most models perform well at higher $k$ values. Llama3 is an exception, showing inconsistent performance and lower scores at higher $k$ values. Similar to the temperature results, larger models do not reliably outperform smaller ones. Specifically, GPT2-XL (1.5B) performs as well as or better than Llama3 (8B), and Qwen2 (7B) outperforms Falcon2 (11B).
For top-$p$ (nucleus) sampling, Figure 24 shows that for all models, performance increases monotonically as $p$ moves from 0.6 to 0.95. The best results occur at $p=0.9$ or $p=0.95$, consistent with previous work by Holtzman et al. (2019). Falcon2 and Qwen2 achieve the highest overall performance, but all models perform well at these higher values. As with the other sampling strategies, larger models do not consistently dominate smaller ones.
It is worth mentioning that combinations of various stochastic decoding strategies are possible, e.g. first adjusting temperature values before truncation with hyperparameters $k$ or $p$. However, this type of analysis lies beyond the scope of the present work.
Contrastive Strategies
Figures 6 through 17 illustrate that optimal Contrastive Search performance depends on balancing the hyperparameters $\alpha$ (degeneration penalty) and $k$ (truncation length). The best overall combination was observed at $\alpha=0.6$ and $k=5$, which aligns with previous work by Su et al. (2022). Other settings of $(\alpha,k)$ such as $(0.8,3)$, $(0.4,20)$, and $(0.4,50)$, also maintained a suitable trade-off between diversity and coherence, performing similar to human references. Very low and very high values of $\alpha$ reduced performance, regardless of the choice of $k$. Specifically, $\alpha=0.2$ produced results similar to beam search, with low diversity and high coherence. In contrast, $\alpha=1.0$ generated outputs with high diversity but very low coherence. These findings align with the work by Su and Xu (2022). Under balanced conditions, all models except Llama3 approached human-level performance. The Llama3 model favored coherence but showed limited diversity, indicating degeneration and suggesting that it may be unsuitable for open-ended generation with this decoding strategy.
We also evaluated the hyperparameter-free Adaptive Contrastive Search. Figure 18 shows that although its MAUVE score is moderate, it achieves a balance of diversity and coherence that produces QText values closer to human references. As indicated in Table 2, it attains the highest diversity (93.95%) and the highest QText score (87.37%) among all tested decoding strategies and hyperparameter combinations. However, its outputs tend to be more coherent than those produced by humans, a pattern observed in most of the examined hyperparameter configurations.
General recommendations
@@
It is important to note that relying solely on automatic evaluation metrics may lead to incomplete assessments of text quality. The results presented here indicate that current automatic metrics do not fully capture human preferences. Previous studies (Garces-Arias et al., 2024) have shown that certain decoding methods, such as CS-based DoubleExp, achieve high scores on metrics like MAUVE but are consistently rejected by human evaluators.
Additionally, stylistic diversity should be considered when choosing decoding strategies. For example, creative writing tasks may benefit from strategies encouraging diversity and moderate coherence, supporting narrative variation. In contrast, tasks focused on factual accuracy and coherence, such as Wikipedia or news generation (as in Wikinews), may require tighter control over coherence. Section A provides detailed performance analyses across different settings, offering further guidance on which strategies yield the best results for each dataset type.
@@ 7 Conclusion
Decoding strategies for large language models are a crucial yet often underexamined aspect of open-ended text generation. This study presents a detailed analysis of how hyperparameter selection across various decoding strategies substantially influences the quality of generated text. The findings highlight the necessity of maintaining a balance between coherence and diversity, as strategies that heavily prioritize one tend to underperform overall. Through extensive sensitivity analysis, we show that the choice of decoding method and its associated hyperparameters can impact text quality as much as, if not more than, model size.
@@
While this study provides insights into how hyperparameters affect the quality of generated text, it also has limitations. First, the impact of our findings on models that have undergone supervised fine-tuning (SFT) remains unexplored. Future work could investigate how these models perform under similar experimental conditions.
Moreover, the analysis was conducted exclusively in English, leaving the generalizability to multilingual settings untested. Previous studies Su and Collier (2023); Garces-Arias et al. (2024) suggest that these findings may generalize to other languages, but further investigation is required.
Finally, this study focuses on analyzing the sensitivity of commonly used automatic metrics, which are inherently influenced by the models employed in the evaluation process. For example, coherence is assessed using the OPT 2.7B model (Zhang et al., 2022), and MAUVE is computed using GPT-2-based embeddings. The potential biases introduced by these model choices have not been fully explored, and addressing this in future research may offer valuable insights.
@@ Acknowledgments
We would like to extend our gratitude to David Garces Arias for his assistance in the Human Evaluation design. Matthias Aßenmacher was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the National Research Data Infrastructure – NFDI 27/1 - 460037581.
Appendix A Appendix
A.1 Case Studies
@@
A.2 Performance per Model, Dataset and Decoding Strategy
Remark: In the following, we exclude Mistralv01 for visualization purposes, as its behavior closely mirrors that of Mistralv03, which employs the same architecture but was trained with additional data.
@@ A.2 Performance per Model, Dataset and Decoding Strategy
Figure 4: Effect of beam width on metric behavior.
@@
A.3 Correlation of Human evaluation and Automatic metrics
Figure 25: Correlation between human judgments and automatic metrics. Human scores show moderate, statistically significant positive correlations with QText and MAUVE. Coherence shows a low, statistically significant positive correlation, while Diversity exhibits a very weak correlation that is not statistically significant.
A.4 Human Evaluation Form